OpenStudy (anonymous):

medal and fan Solve the equation. - 1/3 + w = 5 3/8

DivineSolar (divinesolar):

Let's solve your equation step-by-step. 1 3 +w= (5)(3) 8 Step 1: Simplify both sides of the equation. w+ 1 3 = 15 8 Step 2: Subtract 1/3 from both sides. w+ 1 3 − 1 3 = 15 8 − 1 3 w= 37 24 Answer: w= 37 24

DivineSolar (divinesolar):

Just refresh the page if you see question marks diamonds.

OpenStudy (anonymous):

Thanks

DivineSolar (divinesolar):

Sure thing.

DivineSolar (divinesolar):

Chaneg that answer @Go_to_Sleep forgot to put in the negative! Let's solve your equation step-by-step. −1 3 +w= (5)(3) 8 Step 1: Simplify both sides of the equation. w+ −1 3 = 15 8 Step 2: Add 1/3 to both sides. w+ −1 3 + 1 3 = 15 8 + 1 3 w= 53 24 Answer: w= 53 24

OpenStudy (anonymous):

lol. Thanks again

DivineSolar (divinesolar):

Sorry about that error XD

OpenStudy (anonymous):

It's okay.